Alex, Alex, and Cody are swimming on a balmy 26°C day in water where the speed of sound is 1400 m/s. They have not handed in their labs so Miss Takken sounds like a foghorn to get their attention.  They hear the sound underwater first and hear the sound again in the air 3 seconds later. How far away is Miss Takken from the students?

 




1
Expert's answer
2021-06-14T13:58:55-0400

Find the speed of sound in air at 26°:


"c_\text{air}=331.3\sqrt{1+\frac{T}{273.15}}=\\\space\\
=331.3\sqrt{1+\frac{26}{273.15}}=346.7\text{ m/s}."

The distance to the teacher:


"d=c_\text{air}t=346.7·3=1040.1\text{ m}."